1 The Cyrillic codepages

There are several widely used Cyrillic codepages. Currently, we define here the following codepages:

• cp 866 is the standard MS-DOS Russian codepage. There are also several codepages in use, which are very similar to cp 866. These are: so-called “Cyrillic Alternative codepage” (or Alternative Variant of cp 866), Modified Alternative Variant, New Alternative Variant, and experimental Tatarian codepage. The differences take place in the range 0xf2–0xfe. All these ‘Alternative’ codepages are also supported.

• cp 855 is the standard MS-DOS Cyrillic codepage.

• cp 1251 is the standard MS Windows Cyrillic codepage.

• pt 154 is a Windows Cyrillic Asian codepage developed in ParaType. It is a variant of Windows Cyrillic codepage.

• koi8-r is a standard codepage widely used in UNIX-like systems for Russian language support. It is specified in RFC 1489. The situation with koi8-r is somewhat similar to the one with cp 866: there are also several similar codepages in use, which coincide with koi8-r for all Russian letters, but add some other Cyrillic letters. These codepages include: koi8-u (it is a variant of the koi8-r codepage with some Ukrainian letters added), koi8-ru (it is described in a draft RFC document specifying the widely used character set for mail and news exchange in the Ukrainian internet community as well as for presentating WWW information resources in the Ukrainian language), and ISO-IR-111 ECMA Cyrillic Code Page. All these codepages are supported also.

• ISO 8859-5 Cyrillic codepage (also called ISO-IR-144).

• Apple Macintosh Cyrillic (Microsoft cp 10007) codepage.

• Apple Macintosh Ukrainian codepage (very similar to the previous codepage).

• pt 254 is a Macintosh Cyrillic Asian codepage developed in ParaType. It is a variant of Macintosh Cyrillic codepage.

• Bulgarian MIK (BDS) codepage.

• Mongolian codepages: CTT, DBK, MNK, MOS, NCC, MLS.

For all codepages, one of T2* (or X2) encoding is needed. To access some characters (e.g. \textregistered, \textbrokenbar) present in some codepages, T1 and TS1 are necessary also. However, if the characters used from these codepages will be limited only to Russian letters, it is sufficient to have old LH fonts with LCY or OT2 encoding. In this case, characters which are absent in the font will cause error messages.

Note that the following composite glyphs (using accents) are not ‘named’ here:

\CYRGJE (\'\CYRG), \cyrgje (\'\cyrg), \CYRKJE (\'\CYRK), \cyrkje (\'\cyrk).

Also, \@tabacckludge' is used instead of \' because of the tabbing environment.

1.1 Additional Copyright notice(s)

 1 ⟨CTT | DBK | MNK | MOS | NCC | MLS⟩% (C) Copyright 1999 by Oliver Corff.
 2 ⟨MIK⟩% (C) Copyright 1999 by Georgi Boshnakov, Guentcho Skordev.

1.2 Headers

 3 %\NeedsTeXFormat{LaTeX2e}[1995/12/01]
 4 ⟨cp866 & std⟩\ProvidesFile{cp866.def}
 5 ⟨cp866 & AV⟩\ProvidesFile{cp866av.def}
 6 ⟨cp866 & MAV⟩\ProvidesFile{cp866mav.def}
 7 ⟨cp866 & NAV⟩\ProvidesFile{cp866nav.def}
 8 ⟨cp866 & Tatar⟩\ProvidesFile{cp866tat.def}
 9 ⟨cp1251⟩\ProvidesFile{cp1251.def}
10 ⟨pt154⟩\ProvidesFile{pt154.def}
11 ⟨cp855⟩\ProvidesFile{cp855.def}
12 ⟨koi8 & koi8r⟩\ProvidesFile{koi8-r.def}
13 ⟨koi8 & koi8ru⟩\ProvidesFile{koi8-ru.def}
14 ⟨koi8 & isoir111⟩\ProvidesFile{isoir111.def}
15 ⟨koi8 & koi8u⟩\ProvidesFile{koi8-u.def}
16 ⟨ISO88595⟩\ProvidesFile{iso88595.def}
17 ⟨maccyrillic⟩\ProvidesFile{maccyr.def}
18 ⟨macukrainian⟩\ProvidesFile{macukr.def}
19 ⟨pt254⟩\ProvidesFile{pt254.def}
20 ⟨MIK⟩\ProvidesFile{mik.def}
21 ⟨CTT⟩\ProvidesFile{ctt.def}
22 ⟨DBK⟩\ProvidesFile{dbk.def}
23 ⟨MNK⟩\ProvidesFile{mnk.def}
24 ⟨MOS⟩\ProvidesFile{mos.def}
25 ⟨NCC⟩\ProvidesFile{ncc.def}
26 ⟨MLS⟩\ProvidesFile{mls.def}
27 [1999/11/16 v1.0d Input encoding file]

Insert a \makeatletter at the beginning of all .def files.

28 \makeatletter

29 ⟨∗cp866&!NAV | cp855 | koi8r | koi8ru | MIK⟩

30 \ProvideTextCommandDefault{\textblacksquare}
31 {\vrule \@width .3em \@height .4em \@depth -.1em\relax}

32 ⟨/cp866&!NAV | cp855 | koi8r | koi8ru | MIK⟩ 33 ⟨∗cp866 & std | cp866 & MAV | cp1251 | koi8r | maccyrillic | macukrainian | MIK | pt154 | pt254⟩

34 \ProvideTextCommandDefault{\textdegree}{\ensuremath{{^\circ}}}

35 ⟨/cp866 & std | cp866 & MAV | cp1251 | koi8r | maccyrillic | macukrainian | MIK | pt154 | pt254⟩ 36 ⟨∗cp1251⟩

37 \ProvideTextCommandDefault{\textbrokenbar}
38 {\TextSymbolUnavailable\textbrokenbar}
39 \ProvideTextCommandDefault{\texteuro}
40 {\TextSymbolUnavailable\texteuro}

41 ⟨/cp1251⟩

42 ⟨cp866 & MAV | koi8r | MIK⟩\providecommand{\mathtwosuperior}{{^2}}
43 ⟨cp866 & MAV | MIK⟩\providecommand{\mathnsuperior}{{^n}}

1.3 Microsoft cp 866

44 ⟨∗cp866⟩

45 \DeclareInputText{128}{\CYRA}
46 \DeclareInputText{129}{\CYRB}
47 \DeclareInputText{130}{\CYRV}

48 \DeclareInputText{131}{\CYRG}
49 \DeclareInputText{132}{\CYRD}
50 \DeclareInputText{133}{\CYRE}
51 \DeclareInputText{134}{\CYRZH}
52 \DeclareInputText{135}{\CYRZ}
53 \DeclareInputText{136}{\CYRI}
54 \DeclareInputText{137}{\CYRISHRT}
55 \DeclareInputText{138}{\CYRK}
56 \DeclareInputText{139}{\CYRL}
57 \DeclareInputText{140}{\CYRM}
58 \DeclareInputText{141}{\CYRN}
59 \DeclareInputText{142}{\CYRO}
60 \DeclareInputText{143}{\CYRP}
61 \DeclareInputText{144}{\CYRR}
62 \DeclareInputText{145}{\CYRS}
63 \DeclareInputText{146}{\CYRT}
64 \DeclareInputText{147}{\CYRU}
65 \DeclareInputText{148}{\CYRF}
66 \DeclareInputText{149}{\CYRH}
67 \DeclareInputText{150}{\CYRC}
68 \DeclareInputText{151}{\CYRCH}
69 \DeclareInputText{152}{\CYRSH}
70 \DeclareInputText{153}{\CYRSHCH}
71 \DeclareInputText{154}{\CYRHRDSN}
72 \DeclareInputText{155}{\CYRERY}
73 \DeclareInputText{156}{\CYRSFTSN}
74 \DeclareInputText{157}{\CYREREV}
75 \DeclareInputText{158}{\CYRYU}
76 \DeclareInputText{159}{\CYRYA}

77 %

78 \DeclareInputText{160}{\cyra}
79 \DeclareInputText{161}{\cyrb}
80 \DeclareInputText{162}{\cyrv}
81 \DeclareInputText{163}{\cyrg}
82 \DeclareInputText{164}{\cyrd}
83 \DeclareInputText{165}{\cyre}
84 \DeclareInputText{166}{\cyrzh}
85 \DeclareInputText{167}{\cyrz}
86 \DeclareInputText{168}{\cyri}
87 \DeclareInputText{169}{\cyrishrt}
88 \DeclareInputText{170}{\cyrk}
89 \DeclareInputText{171}{\cyrl}
90 \DeclareInputText{172}{\cyrm}
91 \DeclareInputText{173}{\cyrn}
92 \DeclareInputText{174}{\cyro}
93 \DeclareInputText{175}{\cyrp}
94 \DeclareInputText{224}{\cyrr}
95 \DeclareInputText{225}{\cyrs}
96 \DeclareInputText{226}{\cyrt}
97 \DeclareInputText{227}{\cyru}
98 \DeclareInputText{228}{\cyrf}
99 \DeclareInputText{229}{\cyrh}
100 \DeclareInputText{230}{\cyrc}
101 \DeclareInputText{231}{\cyrch}

102 \DeclareInputText{232}{\cyrsh}
103 \DeclareInputText{233}{\cyrshch}
104 \DeclareInputText{234}{\cyrhrdsn}
105 \DeclareInputText{235}{\cyrery}
106 \DeclareInputText{236}{\cyrsftsn}
107 \DeclareInputText{237}{\cyrerev}
108 \DeclareInputText{238}{\cyryu}
109 \DeclareInputText{239}{\cyrya}

110 %

111 \DeclareInputText{240}{\CYRYO}
112 \DeclareInputText{241}{\cyryo}

The following block corresponds to the standard cp 866 codepage:

113 ⟨∗std⟩

114 \DeclareInputText{242}{\CYRIE}
115 \DeclareInputText{243}{\cyrie}
116 \DeclareInputText{244}{\CYRYI}
117 \DeclareInputText{245}{\cyryi}
118 \DeclareInputText{246}{\CYRUSHRT}
119 \DeclareInputText{247}{\cyrushrt}
120 \DeclareInputText{248}{\textdegree}
121 \DeclareInputText{249}{\textbullet}
122 \DeclareInputText{250}{\textperiodcentered}
123 \DeclareInputMath{251}{\surd}
124 \DeclareInputText{252}{\textnumero}
125 \DeclareInputText{253}{\textcurrency}
126 \DeclareInputText{254}{\textblacksquare}

127 ⟨/std⟩

The following block corresponds to the so called Alternative Variant (AV) of cp 866:

128 ⟨∗AV⟩

129 % 0xf2 LOW ACUTE ACCENT
130 % 0xf3 LOW GRAVE ACCENT
131 % 0xf4 HIGH ACUTE ACCENT
132 % 0xf5 HIGH GRAVE ACCENT
133 \DeclareInputMath{246}{\rightarrow}
134 \DeclareInputMath{247}{\leftarrow}
135 \DeclareInputMath{248}{\downarrow}
136 \DeclareInputMath{249}{\uparrow}
137 \DeclareInputMath{250}{\div}
138 \DeclareInputMath{251}{\pm}
139 \DeclareInputText{252}{\textnumero}
140 \DeclareInputText{253}{\textcurrency}
141 \DeclareInputText{254}{\textblacksquare}

142 ⟨/AV⟩

The following block corresponds to the so called Modified Alternative Variant (MAV) of cp 866. Symbols 0xf2 through 0xfd match standard IBM coding (MS code page 437):

143 ⟨∗MAV⟩

144 \DeclareInputMath{242}{\geq}
145 \DeclareInputMath{243}{\leq}
146 % 0xf4 TOP HALF INTEGRAL
147 % 0xf5 BOTTOM HALF INTEGRAL

148 \DeclareInputMath{246}{\div}
149 \DeclareInputMath{247}{\sim}
150 \DeclareInputText{248}{\textdegree}
151 \DeclareInputText{249}{\textbullet}
152 \DeclareInputText{250}{\textperiodcentered}
153 \DeclareInputMath{251}{\surd}
154 \DeclareInputMath{252}{\mathnsuperior}
155 \DeclareInputMath{253}{\mathtwosuperior}
156 \DeclareInputText{254}{\textblacksquare}

157 ⟨/MAV⟩

The following block corresponds to the yet another modern modification of cp 866:

158 ⟨∗NAV⟩

159 \DeclareInputText{242}{\CYRGUP}
160 \DeclareInputText{243}{\cyrgup}
161 \DeclareInputText{244}{\CYRIE}
162 \DeclareInputText{245}{\cyrie}
163 \DeclareInputText{246}{\CYRII}
164 \DeclareInputText{247}{\cyrii}
165 \DeclareInputText{248}{\CYRYI}
166 \DeclareInputText{249}{\cyryi}
167 \DeclareInputText{250}{\CYRUSHRT}
168 \DeclareInputText{251}{\cyrushrt}
169 \DeclareInputText{252}{\textnumero}
170 % ? left European quotes:
171 \DeclareInputText{253}{\guillemotleft}
172 % ? right European quotes:
173 \DeclareInputText{254}{\guillemotright}

174 ⟨/NAV⟩

The following block corresponds to the experimental Tatarian modification of cp 866. Information was taken from the LH fonts.

175 ⟨∗Tatar⟩

176 \DeclareInputText{242}{\CYRSCHWA}
177 \DeclareInputText{243}{\cyrschwa}
178 \DeclareInputText{244}{\CYROTLD}
179 \DeclareInputText{245}{\cyrotld}
180 \DeclareInputText{246}{\CYRY}
181 \DeclareInputText{247}{\cyry}
182 \DeclareInputText{248}{\CYRZHDSC}
183 \DeclareInputText{249}{\cyrzhdsc}
184 \DeclareInputText{250}{\CYRNDSC}
185 \DeclareInputText{251}{\cyrndsc}
186 \DeclareInputText{252}{\CYRSHHA}
187 \DeclareInputText{253}{\cyrshha}
188 % ? was not explicitly declared:
189 \DeclareInputText{254}{\textblacksquare}

190 ⟨/Tatar⟩

191 \DeclareInputText{255}{\nobreakspace}

192 ⟨/cp866⟩

1.4 Microsoft cp 855

193 ⟨∗cp855⟩

194 \DeclareInputText{128}{\cyrdje}
195 \DeclareInputText{129}{\CYRDJE}
196 \DeclareInputText{130}{\@tabacckludge'\cyrg}
197 \DeclareInputText{131}{\@tabacckludge'\CYRG}
198 \DeclareInputText{132}{\cyryo}
199 \DeclareInputText{133}{\CYRYO}
200 \DeclareInputText{134}{\cyrie}
201 \DeclareInputText{135}{\CYRIE}
202 \DeclareInputText{136}{\cyrdze}
203 \DeclareInputText{137}{\CYRDZE}
204 \DeclareInputText{138}{\cyrii}
205 \DeclareInputText{139}{\CYRII}
206 \DeclareInputText{140}{\cyryi}
207 \DeclareInputText{141}{\CYRYI}
208 \DeclareInputText{142}{\cyrje}
209 \DeclareInputText{143}{\CYRJE}
210 \DeclareInputText{144}{\cyrlje}
211 \DeclareInputText{145}{\CYRLJE}
212 \DeclareInputText{146}{\cyrnje}
213 \DeclareInputText{147}{\CYRNJE}
214 \DeclareInputText{148}{\cyrtshe}
215 \DeclareInputText{149}{\CYRTSHE}
216 \DeclareInputText{150}{\@tabacckludge'\cyrk}
217 \DeclareInputText{151}{\@tabacckludge'\CYRK}
218 \DeclareInputText{152}{\cyrushrt}
219 \DeclareInputText{153}{\CYRUSHRT}
220 \DeclareInputText{154}{\cyrdzhe}
221 \DeclareInputText{155}{\CYRDZHE}
222 \DeclareInputText{156}{\cyryu}
223 \DeclareInputText{157}{\CYRYU}
224 \DeclareInputText{158}{\cyrhrdsn}
225 \DeclareInputText{159}{\CYRHRDSN}
226 \DeclareInputText{160}{\cyra}
227 \DeclareInputText{161}{\CYRA}
228 \DeclareInputText{162}{\cyrb}
229 \DeclareInputText{163}{\CYRB}
230 \DeclareInputText{164}{\cyrc}
231 \DeclareInputText{165}{\CYRC}
232 \DeclareInputText{166}{\cyrd}
233 \DeclareInputText{167}{\CYRD}
234 \DeclareInputText{168}{\cyre}
235 \DeclareInputText{169}{\CYRE}
236 \DeclareInputText{170}{\cyrf}
237 \DeclareInputText{171}{\CYRF}
238 \DeclareInputText{172}{\cyrg}
239 \DeclareInputText{173}{\CYRG}
240 \DeclareInputText{174}{\guillemotleft}
241 \DeclareInputText{175}{\guillemotright}
242 % 0xb0 LIGHT SHADE
243 % 0xb1 MEDIUM SHADE
244 % 0xb2 DARK SHADE
245 % 0xb3 BOX DRAWINGS LIGHT VERTICAL
246 % 0xb4 BOX DRAWINGS LIGHT VERTICAL AND LEFT
247 \DeclareInputText{181}{\cyrh}

248 \DeclareInputText{182}{\CYRH}
249 \DeclareInputText{183}{\cyri}
250 \DeclareInputText{184}{\CYRI}
251 % 0xb9 BOX DRAWINGS DOUBLE VERTICAL AND LEFT
252 % 0xba BOX DRAWINGS DOUBLE VERTICAL
253 % 0xbb BOX DRAWINGS DOUBLE DOWN AND LEFT
254 % 0xbc BOX DRAWINGS DOUBLE UP AND LEFT
255 \DeclareInputText{189}{\cyrishrt}
256 \DeclareInputText{190}{\CYRISHRT}
257 % 0xbf BOX DRAWINGS LIGHT DOWN AND LEFT
258 % 0xc0 BOX DRAWINGS LIGHT UP AND RIGHT
259 % 0xc1 BOX DRAWINGS LIGHT UP AND HORIZONTAL
260 % 0xc2 BOX DRAWINGS LIGHT DOWN AND HORIZONTAL
261 % 0xc3 BOX DRAWINGS LIGHT VERTICAL AND RIGHT
262 % 0xc4 BOX DRAWINGS LIGHT HORIZONTAL
263 % 0xc5 BOX DRAWINGS LIGHT VERTICAL AND HORIZONTAL
264 \DeclareInputText{198}{\cyrk}
265 \DeclareInputText{199}{\CYRK}
266 % 0xc8 BOX DRAWINGS DOUBLE UP AND RIGHT
267 % 0xc9 BOX DRAWINGS DOUBLE DOWN AND RIGHT
268 % 0xca BOX DRAWINGS DOUBLE UP AND HORIZONTAL
269 % 0xcb BOX DRAWINGS DOUBLE DOWN AND HORIZONTAL
270 % 0xcc BOX DRAWINGS DOUBLE VERTICAL AND RIGHT
271 % 0xcd BOX DRAWINGS DOUBLE HORIZONTAL
272 % 0xce BOX DRAWINGS DOUBLE VERTICAL AND HORIZONTAL
273 \DeclareInputText{207}{\textcurrency}
274 \DeclareInputText{208}{\cyrl}
275 \DeclareInputText{209}{\CYRL}
276 \DeclareInputText{210}{\cyrm}
277 \DeclareInputText{211}{\CYRM}
278 \DeclareInputText{212}{\cyrn}
279 \DeclareInputText{213}{\CYRN}
280 \DeclareInputText{214}{\cyro}
281 \DeclareInputText{215}{\CYRO}
282 \DeclareInputText{216}{\cyrp}
283 % 0xd9 BOX DRAWINGS LIGHT UP AND LEFT
284 % 0xda BOX DRAWINGS LIGHT DOWN AND RIGHT
285 % 0xdb FULL BLOCK
286 % 0xdc LOWER HALF BLOCK
287 \DeclareInputText{221}{\CYRP}
288 \DeclareInputText{222}{\cyrya}
289 % 0xdf UPPER HALF BLOCK
290 \DeclareInputText{224}{\CYRYA}
291 \DeclareInputText{225}{\cyrr}
292 \DeclareInputText{226}{\CYRR}
293 \DeclareInputText{227}{\cyrs}
294 \DeclareInputText{228}{\CYRS}
295 \DeclareInputText{229}{\cyrt}
296 \DeclareInputText{230}{\CYRT}
297 \DeclareInputText{231}{\cyru}
298 \DeclareInputText{232}{\CYRU}
299 \DeclareInputText{233}{\cyrzh}
300 \DeclareInputText{234}{\CYRZH}
301 \DeclareInputText{235}{\cyrv}

302 \DeclareInputText{236}{\CYRV}
303 \DeclareInputText{237}{\cyrsftsn}
304 \DeclareInputText{238}{\CYRSFTSN}
305 \DeclareInputText{239}{\textnumero}
306 \DeclareInputText{240}{\-}
307 \DeclareInputText{241}{\cyrery}
308 \DeclareInputText{242}{\CYRERY}
309 \DeclareInputText{243}{\cyrz}
310 \DeclareInputText{244}{\CYRZ}
311 \DeclareInputText{245}{\cyrsh}
312 \DeclareInputText{246}{\CYRSH}
313 \DeclareInputText{247}{\cyrerev}
314 \DeclareInputText{248}{\CYREREV}
315 \DeclareInputText{249}{\cyrshch}
316 \DeclareInputText{250}{\CYRSHCH}
317 \DeclareInputText{251}{\cyrch}
318 \DeclareInputText{252}{\CYRCH}
319 \DeclareInputText{253}{\S}
320 \DeclareInputText{254}{\textblacksquare}
321 \DeclareInputText{255}{\nobreakspace}

322 ⟨/cp855⟩

1.5 Microsoft cp 1251 and ParaType pt 154

323 ⟨∗cp1251 | pt154⟩

324 \DeclareInputText{192}{\CYRA}
325 \DeclareInputText{193}{\CYRB}
326 \DeclareInputText{194}{\CYRV}
327 \DeclareInputText{195}{\CYRG}
328 \DeclareInputText{196}{\CYRD}
329 \DeclareInputText{197}{\CYRE}
330 \DeclareInputText{198}{\CYRZH}
331 \DeclareInputText{199}{\CYRZ}
332 \DeclareInputText{200}{\CYRI}
333 \DeclareInputText{201}{\CYRISHRT}
334 \DeclareInputText{202}{\CYRK}
335 \DeclareInputText{203}{\CYRL}
336 \DeclareInputText{204}{\CYRM}
337 \DeclareInputText{205}{\CYRN}
338 \DeclareInputText{206}{\CYRO}
339 \DeclareInputText{207}{\CYRP}
340 \DeclareInputText{208}{\CYRR}
341 \DeclareInputText{209}{\CYRS}
342 \DeclareInputText{210}{\CYRT}
343 \DeclareInputText{211}{\CYRU}
344 \DeclareInputText{212}{\CYRF}
345 \DeclareInputText{213}{\CYRH}
346 \DeclareInputText{214}{\CYRC}
347 \DeclareInputText{215}{\CYRCH}
348 \DeclareInputText{216}{\CYRSH}
349 \DeclareInputText{217}{\CYRSHCH}
350 \DeclareInputText{218}{\CYRHRDSN}
351 \DeclareInputText{219}{\CYRERY}
352 \DeclareInputText{220}{\CYRSFTSN}
353 \DeclareInputText{221}{\CYREREV}

354 \DeclareInputText{222}{\CYRYU}
355 \DeclareInputText{223}{\CYRYA}

356 %

357 \DeclareInputText{224}{\cyra}
358 \DeclareInputText{225}{\cyrb}
359 \DeclareInputText{226}{\cyrv}
360 \DeclareInputText{227}{\cyrg}
361 \DeclareInputText{228}{\cyrd}
362 \DeclareInputText{229}{\cyre}
363 \DeclareInputText{230}{\cyrzh}
364 \DeclareInputText{231}{\cyrz}
365 \DeclareInputText{232}{\cyri}
366 \DeclareInputText{233}{\cyrishrt}
367 \DeclareInputText{234}{\cyrk}
368 \DeclareInputText{235}{\cyrl}
369 \DeclareInputText{236}{\cyrm}
370 \DeclareInputText{237}{\cyrn}
371 \DeclareInputText{238}{\cyro}
372 \DeclareInputText{239}{\cyrp}
373 \DeclareInputText{240}{\cyrr}
374 \DeclareInputText{241}{\cyrs}
375 \DeclareInputText{242}{\cyrt}
376 \DeclareInputText{243}{\cyru}
377 \DeclareInputText{244}{\cyrf}
378 \DeclareInputText{245}{\cyrh}
379 \DeclareInputText{246}{\cyrc}
380 \DeclareInputText{247}{\cyrch}
381 \DeclareInputText{248}{\cyrsh}
382 \DeclareInputText{249}{\cyrshch}
383 \DeclareInputText{250}{\cyrhrdsn}
384 \DeclareInputText{251}{\cyrery}
385 \DeclareInputText{252}{\cyrsftsn}
386 \DeclareInputText{253}{\cyrerev}
387 \DeclareInputText{254}{\cyryu}
388 \DeclareInputText{255}{\cyrya}

389 %

390 ⟨cp1251⟩\DeclareInputText{128}{\CYRDJE}
391 ⟨cp1251⟩\DeclareInputText{129}{\@tabacckludge'\CYRG}
392 ⟨pt154⟩\DeclareInputText{128}{\CYRZHDSC}
393 ⟨pt154⟩\DeclareInputText{129}{\CYRGHCRS}
394 \DeclareInputText{130}{\quotesinglbase}
395 ⟨cp1251⟩\DeclareInputText{131}{\@tabacckludge'\cyrg}
396 ⟨pt154⟩\DeclareInputText{131}{\cyrghcrs}
397 \DeclareInputText{132}{\quotedblbase}
398 \DeclareInputText{133}{\dots}
399 \DeclareInputText{134}{\dag}

400 ⟨∗cp1251⟩

401 \DeclareInputText{135}{\ddag}
402 \DeclareInputText{136}{\texteuro}
403 \DeclareInputText{137}{\textperthousand}
404 \DeclareInputText{138}{\CYRLJE}

405 ⟨/cp1251⟩ 406 ⟨∗pt154⟩

407 \DeclareInputText{135}{\CYRY}

408 \DeclareInputText{136}{\CYRHDSC}
409 \DeclareInputText{137}{\cyry}
410 \DeclareInputText{138}{\CYRKBEAK}

411 ⟨/pt154⟩

412 \DeclareInputText{139}{\guilsinglleft}

413 ⟨∗cp1251⟩

414 \DeclareInputText{140}{\CYRNJE}
415 \DeclareInputText{141}{\@tabacckludge'\CYRK}
416 \DeclareInputText{142}{\CYRTSHE}
417 \DeclareInputText{143}{\CYRDZHE}
418 \DeclareInputText{144}{\cyrdje}

419 ⟨/cp1251⟩ 420 ⟨∗pt154⟩

421 \DeclareInputText{140}{\CYRNDSC}
422 \DeclareInputText{141}{\CYRKDSC}
423 \DeclareInputText{142}{\CYRSHHA}
424 \DeclareInputText{143}{\CYRCHVCRS}
425 \DeclareInputText{144}{\cyrzhdsc}

426 ⟨/pt154⟩

427 \DeclareInputText{145}{\textquoteleft}
428 \DeclareInputText{146}{\textquoteright}
429 \DeclareInputText{147}{\textquotedblleft}
430 \DeclareInputText{148}{\textquotedblright}
431 \DeclareInputText{149}{\textbullet}
432 \DeclareInputText{150}{\textendash}
433 \DeclareInputText{151}{\textemdash}
434 % 0x98 undefined in cp1251
435 ⟨pt154⟩\DeclareInputText{152}{\cyrhdsc}
436 \DeclareInputText{153}{\texttrademark}
437 ⟨cp1251⟩\DeclareInputText{154}{\cyrlje}
438 ⟨pt154⟩\DeclareInputText{154}{\cyrkbeak}
439 \DeclareInputText{155}{\guilsinglright}

440 ⟨∗cp1251⟩

441 \DeclareInputText{156}{\cyrnje}
442 \DeclareInputText{157}{\@tabacckludge'\cyrk}
443 \DeclareInputText{158}{\cyrtshe}
444 \DeclareInputText{159}{\cyrdzhe}

445 ⟨/cp1251⟩ 446 ⟨∗pt154⟩

447 \DeclareInputText{156}{\cyrndsc}
448 \DeclareInputText{157}{\cyrkdsc}
449 \DeclareInputText{158}{\cyrshha}
450 \DeclareInputText{159}{\cyrchvcrs}

451 ⟨/pt154⟩

452 \DeclareInputText{160}{\nobreakspace}
453 \DeclareInputText{161}{\CYRUSHRT}
454 \DeclareInputText{162}{\cyrushrt}
455 \DeclareInputText{163}{\CYRJE}

456 ⟨∗cp1251⟩

457 \DeclareInputText{164}{\textcurrency}
458 \DeclareInputText{165}{\CYRGUP}
459 \DeclareInputText{166}{\textbrokenbar}

460 ⟨/cp1251⟩ 461 ⟨∗pt154⟩

462 \DeclareInputText{164}{\CYROTLD}
463 \DeclareInputText{165}{\CYRZDSC}
464 \DeclareInputText{166}{\CYRYHCRS}

465 ⟨/pt154⟩

466 \DeclareInputText{167}{\S}
467 \DeclareInputText{168}{\CYRYO}
468 \DeclareInputText{169}{\copyright}
469 ⟨cp1251⟩\DeclareInputText{170}{\CYRIE}
470 ⟨pt154⟩\DeclareInputText{170}{\CYRSCHWA}
471 \DeclareInputText{171}{\guillemotleft}
472 \DeclareInputMath{172}{\lnot}
473 \DeclareInputText{173}{\-}
474 \DeclareInputText{174}{\textregistered}
475 ⟨cp1251⟩\DeclareInputText{175}{\CYRYI}
476 ⟨pt154⟩\DeclareInputText{175}{\CYRKVCRS}
477 \DeclareInputText{176}{\textdegree}
478 ⟨cp1251⟩\DeclareInputMath{177}{\pm}
479 ⟨pt154⟩\DeclareInputText{177}{\cyryhcrs}
480 \DeclareInputText{178}{\CYRII}
481 \DeclareInputText{179}{\cyrii}
482 ⟨cp1251⟩\DeclareInputText{180}{\cyrgup}
483 ⟨cp1251⟩\DeclareInputMath{181}{\mu}
484 ⟨pt154⟩\DeclareInputText{180}{\cyrzdsc}
485 ⟨pt154⟩\DeclareInputText{181}{\cyrotld}
486 \DeclareInputText{182}{\P}
487 \DeclareInputText{183}{\textperiodcentered}
488 \DeclareInputText{184}{\cyryo}
489 \DeclareInputText{185}{\textnumero}
490 ⟨cp1251⟩\DeclareInputText{186}{\cyrie}
491 ⟨pt154⟩\DeclareInputText{186}{\cyrschwa}
492 \DeclareInputText{187}{\guillemotright}
493 \DeclareInputText{188}{\cyrje}

494 ⟨∗cp1251⟩

495 \DeclareInputText{189}{\CYRDZE}
496 \DeclareInputText{190}{\cyrdze}
497 \DeclareInputText{191}{\cyryi}

498 ⟨/cp1251⟩ 499 ⟨∗pt154⟩

500 \DeclareInputText{189}{\CYRSDSC}
501 \DeclareInputText{190}{\cyrsdsc}
502 \DeclareInputText{191}{\cyrkvcrs}

503 ⟨/pt154⟩ 504 ⟨/cp1251 | pt154⟩

1.6 The koi8 codepage

505 ⟨∗koi8⟩

506 \DeclareInputText{225}{\CYRA}
507 \DeclareInputText{226}{\CYRB}
508 \DeclareInputText{247}{\CYRV}
509 \DeclareInputText{231}{\CYRG}
510 \DeclareInputText{228}{\CYRD}
511 \DeclareInputText{229}{\CYRE}
512 \DeclareInputText{179}{\CYRYO}
513 \DeclareInputText{246}{\CYRZH}

514 \DeclareInputText{250}{\CYRZ}
515 \DeclareInputText{233}{\CYRI}
516 \DeclareInputText{234}{\CYRISHRT}
517 \DeclareInputText{235}{\CYRK}
518 \DeclareInputText{236}{\CYRL}
519 \DeclareInputText{237}{\CYRM}
520 \DeclareInputText{238}{\CYRN}
521 \DeclareInputText{239}{\CYRO}
522 \DeclareInputText{240}{\CYRP}
523 \DeclareInputText{242}{\CYRR}
524 \DeclareInputText{243}{\CYRS}
525 \DeclareInputText{244}{\CYRT}
526 \DeclareInputText{245}{\CYRU}
527 \DeclareInputText{230}{\CYRF}
528 \DeclareInputText{232}{\CYRH}
529 \DeclareInputText{227}{\CYRC}
530 \DeclareInputText{254}{\CYRCH}
531 \DeclareInputText{251}{\CYRSH}
532 \DeclareInputText{253}{\CYRSHCH}
533 \DeclareInputText{255}{\CYRHRDSN}
534 \DeclareInputText{249}{\CYRERY}
535 \DeclareInputText{248}{\CYRSFTSN}
536 \DeclareInputText{252}{\CYREREV}
537 \DeclareInputText{224}{\CYRYU}
538 \DeclareInputText{241}{\CYRYA}

539 %

540 \DeclareInputText{193}{\cyra}
541 \DeclareInputText{194}{\cyrb}
542 \DeclareInputText{215}{\cyrv}
543 \DeclareInputText{199}{\cyrg}
544 \DeclareInputText{196}{\cyrd}
545 \DeclareInputText{197}{\cyre}
546 \DeclareInputText{163}{\cyryo}
547 \DeclareInputText{214}{\cyrzh}
548 \DeclareInputText{218}{\cyrz}
549 \DeclareInputText{201}{\cyri}
550 \DeclareInputText{202}{\cyrishrt}
551 \DeclareInputText{203}{\cyrk}
552 \DeclareInputText{204}{\cyrl}
553 \DeclareInputText{205}{\cyrm}
554 \DeclareInputText{206}{\cyrn}
555 \DeclareInputText{207}{\cyro}
556 \DeclareInputText{208}{\cyrp}
557 \DeclareInputText{210}{\cyrr}
558 \DeclareInputText{211}{\cyrs}
559 \DeclareInputText{212}{\cyrt}
560 \DeclareInputText{213}{\cyru}
561 \DeclareInputText{198}{\cyrf}
562 \DeclareInputText{200}{\cyrh}
563 \DeclareInputText{195}{\cyrc}
564 \DeclareInputText{222}{\cyrch}
565 \DeclareInputText{219}{\cyrsh}
566 \DeclareInputText{221}{\cyrshch}
567 \DeclareInputText{223}{\cyrhrdsn}

568 \DeclareInputText{217}{\cyrery}
569 \DeclareInputText{216}{\cyrsftsn}
570 \DeclareInputText{220}{\cyrerev}
571 \DeclareInputText{192}{\cyryu}
572 \DeclareInputText{209}{\cyrya}

1.6.1 koi8-r and relatives (koi8-ru, koi8-u, ISO-IR-111)

0x80–0x9f are unused in the ISO IR-111 Cyrillic Code Page

573 ⟨∗koi8r | koi8ru⟩

574 % 0x80 FORMS LIGHT HORIZONTAL
575 % 0x81 FORMS LIGHT VERTICAL
576 % 0x82 FORMS LIGHT DOWN AND RIGHT
577 % 0x83 FORMS LIGHT DOWN AND LEFT
578 % 0x84 FORMS LIGHT UP AND RIGHT
579 % 0x85 FORMS LIGHT UP AND LEFT
580 % 0x86 FORMS LIGHT VERTICAL AND RIGHT
581 % 0x87 FORMS LIGHT VERTICAL AND LEFT
582 % 0x88 FORMS LIGHT DOWN AND HORIZONTAL
583 % 0x89 FORMS LIGHT UP AND HORIZONTAL
584 % 0x8A FORMS LIGHT VERTICAL AND HORIZONTAL
585 % 0x8B UPPER HALF BLOCK
586 % 0x8C LOWER HALF BLOCK
587 % 0x8D FULL BLOCK
588 % 0x8E LEFT HALF BLOCK
589 % 0x8F RIGHT HALF BLOCK
590 % 0x90 LIGHT SHADE
591 % 0x91 MEDIUM SHADE
592 % 0x92 DARK SHADE

593 ⟨/koi8r | koi8ru⟩ 594 ⟨∗koi8r⟩

595 % 0x93 TOP HALF INTEGRAL

596 ⟨/koi8r⟩

597 ⟨koi8ru⟩\DeclareInputText{147}{\textquotedblleft}

598 ⟨∗koi8r | koi8ru⟩

599 \DeclareInputText{148}{\textblacksquare}
600 \DeclareInputText{149}{\textbullet}

601 ⟨/koi8r | koi8ru⟩ 602 ⟨∗koi8r⟩

603 \DeclareInputMath{150}{\surd}
604 \DeclareInputMath{151}{\sim}
605 \DeclareInputMath{152}{\leq}
606 \DeclareInputMath{153}{\geq}

607 ⟨/koi8r⟩ 608 ⟨∗koi8ru⟩

609 \DeclareInputText{150}{\textquotedblright}
610 \DeclareInputText{151}{\textemdash}
611 \DeclareInputText{152}{\textnumero}
612 \DeclareInputText{153}{\texttrademark}

613 ⟨/koi8ru⟩

614 ⟨koi8r | koi8ru⟩\DeclareInputText{154}{\nobreakspace}

615 ⟨∗koi8r⟩

616 % 0x9B BOTTOM HALF INTEGRAL
617 \DeclareInputText{156}{\textdegree}
618 \DeclareInputMath{157}{\mathtwosuperior}

619 ⟨/koi8r⟩ 620 ⟨∗koi8ru⟩

621 \DeclareInputText{155}{\guillemotright}
622 \DeclareInputText{156}{\textregistered}
623 \DeclareInputText{157}{\guillemotleft}

624 ⟨/koi8ru⟩

625 ⟨koi8r | koi8ru⟩\DeclareInputText{158}{\textperiodcentered}
626 ⟨koi8r⟩\DeclareInputMath{159}{\div}
627 ⟨koi8ru⟩\DeclareInputText{159}{\textcurrency}

628 ⟨∗koi8r | koi8ru⟩

629 % 0xA0 FORMS DOUBLE HORIZONTAL
630 % 0xA1 FORMS DOUBLE VERTICAL
631 % 0xA2 FORMS DOWN SINGLE AND RIGHT DOUBLE

632 ⟨/koi8r | koi8ru⟩ 633 ⟨∗isoir111⟩

634 \DeclareInputText{160}{\nobreakspace}
635 \DeclareInputText{161}{\cyrdje}
636 \DeclareInputText{162}{\@tabacckludge'\cyrg}

637 ⟨/isoir111⟩ 638 ⟨∗koi8r⟩

639 % 0xA4 FORMS DOWN DOUBLE AND RIGHT SINGLE

640 ⟨/koi8r⟩

641 ⟨koi8ru | isoir111 | koi8u⟩\DeclareInputText{164}{\cyrie}

642 ⟨∗koi8r | koi8ru⟩

643 % 0xA5 FORMS DOUBLE DOWN AND RIGHT

644 ⟨/koi8r | koi8ru⟩

645 ⟨isoir111⟩\DeclareInputText{165}{\cyrdze}

646 ⟨∗koi8r⟩

647 % 0xA6 FORMS DOWN SINGLE AND LEFT DOUBLE
648 % 0xA7 FORMS DOWN DOUBLE AND LEFT SINGLE

649 ⟨/koi8r⟩ 650 ⟨∗koi8ru | isoir111 | koi8u⟩

651 \DeclareInputText{166}{\cyrii}
652 \DeclareInputText{167}{\cyryi}

653 ⟨/koi8ru | isoir111 | koi8u⟩ 654 ⟨∗koi8r | koi8ru⟩

655 % 0xA8 FORMS DOUBLE DOWN AND LEFT
656 % 0xA9 FORMS UP SINGLE AND RIGHT DOUBLE
657 % 0xAA FORMS UP DOUBLE AND RIGHT SINGLE
658 % 0xAB FORMS DOUBLE UP AND RIGHT
659 % 0xAC FORMS UP SINGLE AND LEFT DOUBLE

660 ⟨/koi8r | koi8ru⟩ 661 ⟨∗isoir111⟩

662 \DeclareInputText{168}{\cyrje}
663 \DeclareInputText{169}{\cyrlje}
664 \DeclareInputText{170}{\cyrnje}
665 \DeclareInputText{171}{\cyrtshe}
666 \DeclareInputText{172}{\@tabacckludge'\cyrk}

667 ⟨/isoir111⟩ 668 ⟨∗koi8r⟩

669 % 0xAD FORMS UP DOUBLE AND LEFT SINGLE
670 % 0xAE FORMS DOUBLE UP AND LEFT

671 ⟨/koi8r⟩

672 ⟨koi8ru | koi8u⟩\DeclareInputText{173}{\cyrgup}

673 ⟨isoir111⟩\DeclareInputText{173}{\-}
674 ⟨koi8ru | isoir111⟩\DeclareInputText{174}{\cyrushrt}

675 ⟨∗koi8r | koi8ru⟩

676 % 0xAF FORMS VERTICAL SINGLE AND RIGHT DOUBLE
677 % 0xB0 FORMS VERTICAL DOUBLE AND RIGHT SINGLE
678 % 0xB1 FORMS DOUBLE VERTICAL AND RIGHT
679 % 0xB2 FORMS VERTICAL SINGLE AND LEFT DOUBLE

680 ⟨/koi8r | koi8ru⟩ 681 ⟨∗isoir111⟩

682 \DeclareInputText{175}{\cyrdzhe}
683 \DeclareInputText{176}{\textnumero}
684 \DeclareInputText{177}{\CYRDJE}
685 \DeclareInputText{178}{\@tabacckludge'\CYRG}

686 ⟨/isoir111⟩ 687 ⟨∗koi8r⟩

688 % 0xB4 FORMS VERTICAL DOUBLE AND LEFT SINGLE

689 ⟨/koi8r⟩

690 ⟨koi8ru | isoir111 | koi8u⟩\DeclareInputText{180}{\CYRIE}

691 ⟨∗koi8r | koi8ru⟩

692 % 0xB5 FORMS DOUBLE VERTICAL AND LEFT

693 ⟨/koi8r | koi8ru⟩

694 ⟨isoir111⟩\DeclareInputText{181}{\CYRDZE}

695 ⟨∗koi8r⟩

696 % 0xB6 FORMS DOWN SINGLE AND HORIZONTAL DOUBLE
697 % 0xB7 FORMS DOWN DOUBLE AND HORIZONTAL SINGLE

698 ⟨/koi8r⟩ 699 ⟨∗koi8ru | isoir111 | koi8u⟩

700 \DeclareInputText{182}{\CYRII}
701 \DeclareInputText{183}{\CYRYI}

702 ⟨/koi8ru | isoir111 | koi8u⟩ 703 ⟨∗koi8r | koi8ru⟩

704 % 0xB8 FORMS DOUBLE DOWN AND HORIZONTAL
705 % 0xB9 FORMS UP SINGLE AND HORIZONTAL DOUBLE
706 % 0xBA FORMS UP DOUBLE AND HORIZONTAL SINGLE
707 % 0xBB FORMS DOUBLE UP AND HORIZONTAL
708 % 0xBC FORMS VERTICAL SINGLE AND HORIZONTAL DOUBLE

709 ⟨/koi8r | koi8ru⟩ 710 ⟨∗isoir111⟩

711 \DeclareInputText{184}{\CYRJE}
712 \DeclareInputText{185}{\CYRLJE}
713 \DeclareInputText{186}{\CYRNJE}
714 \DeclareInputText{187}{\CYRTSHE}
715 \DeclareInputText{188}{\@tabacckludge'\CYRK}

716 ⟨/isoir111⟩ 717 ⟨∗koi8r⟩

718 % 0xBD FORMS VERTICAL DOUBLE AND HORIZONTAL SINGLE
719 % 0xBE FORMS DOUBLE VERTICAL AND HORIZONTAL

720 ⟨/koi8r⟩

721 ⟨koi8ru | koi8u⟩\DeclareInputText{189}{\CYRGUP}
722 ⟨isoir111⟩\DeclareInputText{189}{\textcurrency}
723 ⟨koi8ru | isoir111⟩\DeclareInputText{190}{\CYRUSHRT}
724 ⟨koi8r | koi8ru⟩\DeclareInputText{191}{\copyright}
725 ⟨isoir111⟩\DeclareInputText{191}{\CYRDZHE}

726 ⟨/koi8⟩

1.7 ISO 8859-5

727 ⟨∗ISO88595⟩

728 \DeclareInputText{160}{\nobreakspace}
729 \DeclareInputText{161}{\CYRYO}
730 \DeclareInputText{162}{\CYRDJE}
731 \DeclareInputText{163}{\@tabacckludge'\CYRG}
732 \DeclareInputText{164}{\CYRIE}
733 \DeclareInputText{165}{\CYRDZE}
734 \DeclareInputText{166}{\CYRII}
735 \DeclareInputText{167}{\CYRYI}
736 \DeclareInputText{168}{\CYRJE}
737 \DeclareInputText{169}{\CYRLJE}
738 \DeclareInputText{170}{\CYRNJE}
739 \DeclareInputText{171}{\CYRTSHE}
740 \DeclareInputText{172}{\@tabacckludge'\CYRK}
741 \DeclareInputText{173}{\-}
742 \DeclareInputText{174}{\CYRUSHRT}
743 \DeclareInputText{175}{\CYRDZHE}

744 %

745 \DeclareInputText{176}{\CYRA}
746 \DeclareInputText{177}{\CYRB}
747 \DeclareInputText{178}{\CYRV}
748 \DeclareInputText{179}{\CYRG}
749 \DeclareInputText{180}{\CYRD}
750 \DeclareInputText{181}{\CYRE}
751 \DeclareInputText{182}{\CYRZH}
752 \DeclareInputText{183}{\CYRZ}
753 \DeclareInputText{184}{\CYRI}
754 \DeclareInputText{185}{\CYRISHRT}
755 \DeclareInputText{186}{\CYRK}
756 \DeclareInputText{187}{\CYRL}
757 \DeclareInputText{188}{\CYRM}
758 \DeclareInputText{189}{\CYRN}
759 \DeclareInputText{190}{\CYRO}
760 \DeclareInputText{191}{\CYRP}
761 \DeclareInputText{192}{\CYRR}
762 \DeclareInputText{193}{\CYRS}
763 \DeclareInputText{194}{\CYRT}
764 \DeclareInputText{195}{\CYRU}
765 \DeclareInputText{196}{\CYRF}
766 \DeclareInputText{197}{\CYRH}
767 \DeclareInputText{198}{\CYRC}
768 \DeclareInputText{199}{\CYRCH}
769 \DeclareInputText{200}{\CYRSH}
770 \DeclareInputText{201}{\CYRSHCH}
771 \DeclareInputText{202}{\CYRHRDSN}
772 \DeclareInputText{203}{\CYRERY}
773 \DeclareInputText{204}{\CYRSFTSN}
774 \DeclareInputText{205}{\CYREREV}
775 \DeclareInputText{206}{\CYRYU}
776 \DeclareInputText{207}{\CYRYA}

777 %

778 \DeclareInputText{208}{\cyra}

779 \DeclareInputText{209}{\cyrb}
780 \DeclareInputText{210}{\cyrv}
781 \DeclareInputText{211}{\cyrg}
782 \DeclareInputText{212}{\cyrd}
783 \DeclareInputText{213}{\cyre}
784 \DeclareInputText{214}{\cyrzh}
785 \DeclareInputText{215}{\cyrz}
786 \DeclareInputText{216}{\cyri}
787 \DeclareInputText{217}{\cyrishrt}
788 \DeclareInputText{218}{\cyrk}
789 \DeclareInputText{219}{\cyrl}
790 \DeclareInputText{220}{\cyrm}
791 \DeclareInputText{221}{\cyrn}
792 \DeclareInputText{222}{\cyro}
793 \DeclareInputText{223}{\cyrp}
794 \DeclareInputText{224}{\cyrr}
795 \DeclareInputText{225}{\cyrs}
796 \DeclareInputText{226}{\cyrt}
797 \DeclareInputText{227}{\cyru}
798 \DeclareInputText{228}{\cyrf}
799 \DeclareInputText{229}{\cyrh}
800 \DeclareInputText{230}{\cyrc}
801 \DeclareInputText{231}{\cyrch}
802 \DeclareInputText{232}{\cyrsh}
803 \DeclareInputText{233}{\cyrshch}
804 \DeclareInputText{234}{\cyrhrdsn}
805 \DeclareInputText{235}{\cyrery}
806 \DeclareInputText{236}{\cyrsftsn}
807 \DeclareInputText{237}{\cyrerev}
808 \DeclareInputText{238}{\cyryu}
809 \DeclareInputText{239}{\cyrya}

810 %

811 \DeclareInputText{240}{\textnumero}
812 \DeclareInputText{241}{\cyryo}
813 \DeclareInputText{242}{\cyrdje}
814 \DeclareInputText{243}{\@tabacckludge'\cyrg}
815 \DeclareInputText{244}{\cyrie}
816 \DeclareInputText{245}{\cyrdze}
817 \DeclareInputText{246}{\cyrii}
818 \DeclareInputText{247}{\cyryi}
819 \DeclareInputText{248}{\cyrje}
820 \DeclareInputText{249}{\cyrlje}
821 \DeclareInputText{250}{\cyrnje}
822 \DeclareInputText{251}{\cyrtshe}
823 \DeclareInputText{252}{\@tabacckludge'\cyrk}
824 \DeclareInputText{253}{\S}
825 \DeclareInputText{254}{\cyrushrt}
826 \DeclareInputText{255}{\cyrdzhe}

827 ⟨/ISO88595⟩

1.8 Apple Macintosh Cyrillic encodings and ParaType pt 254

The MacOS Cyrillic encoding (Microsoft cp 10007) includes the full Cyrillic letter repertory of ISO 8859-5 (although not at the same code points). This covers most of the Slavic languages written with the Cyrillic script.

The MacOS Cyrillic encoding also includes a number of characters needed for the MacOS user interface (e.g. ellipsis, bullet for echoing passwords, copyright sign, etc). All of the characters in MacOS Cyrillic that are also in the MacOS Roman encoding are at the same code points as specified in MacOS Roman. This improves application compatibility (since some naughty applications hard-code the MacOS Roman code points of certain characters).

A variant of MacOS Cyrillic is used for Ukrainian. This character encoding adds upper and lower GHE WITH UPTURN, for a grand total of 2 code point differences from standard MacOS Cyrillic.

828 ⟨∗maccyrillic | macukrainian | pt254⟩

829 \DeclareInputText{128}{\CYRA}
830 \DeclareInputText{129}{\CYRB}
831 \DeclareInputText{130}{\CYRV}
832 \DeclareInputText{131}{\CYRG}
833 \DeclareInputText{132}{\CYRD}
834 \DeclareInputText{133}{\CYRE}
835 \DeclareInputText{134}{\CYRZH}
836 \DeclareInputText{135}{\CYRZ}
837 \DeclareInputText{136}{\CYRI}
838 \DeclareInputText{137}{\CYRISHRT}
839 \DeclareInputText{138}{\CYRK}
840 \DeclareInputText{139}{\CYRL}
841 \DeclareInputText{140}{\CYRM}
842 \DeclareInputText{141}{\CYRN}
843 \DeclareInputText{142}{\CYRO}
844 \DeclareInputText{143}{\CYRP}
845 \DeclareInputText{144}{\CYRR}
846 \DeclareInputText{145}{\CYRS}
847 \DeclareInputText{146}{\CYRT}
848 \DeclareInputText{147}{\CYRU}
849 \DeclareInputText{148}{\CYRF}
850 \DeclareInputText{149}{\CYRH}
851 \DeclareInputText{150}{\CYRC}
852 \DeclareInputText{151}{\CYRCH}
853 \DeclareInputText{152}{\CYRSH}
854 \DeclareInputText{153}{\CYRSHCH}
855 \DeclareInputText{154}{\CYRHRDSN}
856 \DeclareInputText{155}{\CYRERY}
857 \DeclareInputText{156}{\CYRSFTSN}
858 \DeclareInputText{157}{\CYREREV}
859 \DeclareInputText{158}{\CYRYU}
860 \DeclareInputText{159}{\CYRYA}

861 %

862 \DeclareInputText{160}{\dag}
863 \DeclareInputText{161}{\textdegree}
864 ⟨maccyrillic | pt254⟩\DeclareInputText{162}{\textcent}
865 ⟨macukrainian⟩\DeclareInputText{162}{\CYRGUP}
866 \DeclareInputText{163}{\pounds}
867 \DeclareInputText{164}{\S}
868 \DeclareInputText{165}{\textbullet}
869 \DeclareInputText{166}{\P}
870 \DeclareInputText{167}{\CYRII}
871 \DeclareInputText{168}{\textregistered}

872 \DeclareInputText{169}{\copyright}
873 \DeclareInputText{170}{\texttrademark}

874 ⟨∗maccyrillic | macukrainian⟩

875 \DeclareInputText{171}{\CYRDJE}
876 \DeclareInputText{172}{\cyrdje}
877 \DeclareInputMath{173}{\neq}
878 \DeclareInputText{174}{\@tabacckludge'\CYRG}
879 \DeclareInputText{175}{\@tabacckludge'\cyrg}
880 \DeclareInputMath{176}{\infty}
881 \DeclareInputMath{177}{\pm}

882 ⟨/maccyrillic | macukrainian⟩ 883 ⟨∗pt254⟩

884 \DeclareInputText{171}{\CYRZHDSC}
885 \DeclareInputText{172}{\cyrzhdsc}
886 \DeclareInputText{173}{\cyrii}
887 \DeclareInputText{174}{\CYRGHCRS}
888 \DeclareInputText{175}{\cyrghcrs}
889 \DeclareInputText{176}{\CYRZDSC}
890 \DeclareInputText{177}{\cyrzdsc}

891 ⟨/pt254⟩

892 \DeclareInputMath{178}{\leq}
893 \DeclareInputMath{179}{\geq}
894 ⟨maccyrillic | macukrainian⟩\DeclareInputText{180}{\cyrii}

895 ⟨maccyrillic | macukrainian⟩\DeclareInputMath{181}{\mu}

896 ⟨pt254⟩\DeclareInputText{180}{\CYRYHCRS}
897 ⟨pt254⟩\DeclareInputText{181}{\cyrotld}
898 ⟨maccyrillic⟩\DeclareInputMath{182}{\partial}
899 ⟨macukrainian⟩\DeclareInputText{182}{\cyrgup}
900 ⟨pt254⟩\DeclareInputText{182}{\CYRY}

901 %

902 \DeclareInputText{183}{\CYRJE}

903 ⟨∗maccyrillic | macukrainian⟩

904 \DeclareInputText{184}{\CYRIE}
905 \DeclareInputText{185}{\cyrie}
906 \DeclareInputText{186}{\CYRYI}
907 \DeclareInputText{187}{\cyryi}
908 \DeclareInputText{188}{\CYRLJE}
909 \DeclareInputText{189}{\cyrlje}
910 \DeclareInputText{190}{\CYRNJE}
911 \DeclareInputText{191}{\cyrnje}

912 ⟨/maccyrillic | macukrainian⟩ 913 ⟨∗pt254⟩

914 \DeclareInputText{184}{\CYRSCHWA}
915 \DeclareInputText{185}{\cyrschwa}
916 \DeclareInputText{186}{\CYRKVCRS}
917 \DeclareInputText{187}{\cyrkvcrs}
918 \DeclareInputText{188}{\CYRKBEAK}
919 \DeclareInputText{189}{\cyrkbeak}
920 \DeclareInputText{190}{\CYRNDSC}
921 \DeclareInputText{191}{\cyrndsc}

922 ⟨/pt254⟩

923 \DeclareInputText{192}{\cyrje}
924 ⟨maccyrillic | macukrainian⟩\DeclareInputText{193}{\CYRDZE}
925 ⟨pt254⟩\DeclareInputText{193}{\CYRSDSC}

927 \DeclareInputMath{194}{\lnot}

928 ⟨∗maccyrillic | macukrainian⟩

929 \DeclareInputMath{195}{\surd}
930 \DeclareInputText{196}{\textflorin}
931 \DeclareInputMath{197}{\approx}
932 % INCREMENT:
933 \DeclareInputMath{198}{\Delta}

934 ⟨/maccyrillic | macukrainian⟩ 935 ⟨∗pt254⟩

936 \DeclareInputText{195}{\CYRHDSC}
937 \DeclareInputText{196}{\cyryhcrs}
938 \DeclareInputText{197}{\cyrhdsc}
939 \DeclareInputText{198}{\cyry}

940 ⟨/pt254⟩

941 \DeclareInputText{199}{\guillemotleft}
942 \DeclareInputText{200}{\guillemotright}
943 % HORIZONTAL ELLIPSIS:
944 \DeclareInputText{201}{\dots}
945 \DeclareInputText{202}{\nobreakspace}

946 % 947 ⟨∗maccyrillic | macukrainian⟩

948 \DeclareInputText{203}{\CYRTSHE}
949 \DeclareInputText{204}{\cyrtshe}
950 \DeclareInputText{205}{\@tabacckludge'\CYRK}
951 \DeclareInputText{206}{\@tabacckludge'\cyrk}
952 \DeclareInputText{207}{\cyrdze}

953 ⟨/maccyrillic | macukrainian⟩ 954 ⟨∗pt254⟩

955 \DeclareInputText{203}{\CYRSHHA}
956 \DeclareInputText{204}{\cyrshha}
957 \DeclareInputText{205}{\CYRKDSC}
958 \DeclareInputText{206}{\cyrkdsc}
959 \DeclareInputText{207}{\cyrsdsc}

960 ⟨/pt254⟩ 961 %

962 \DeclareInputText{208}{\textendash}
963 \DeclareInputText{209}{\textemdash}
964 \DeclareInputText{210}{\textquotedblleft}
965 \DeclareInputText{211}{\textquotedblright}
966 \DeclareInputText{212}{\textquoteleft}
967 \DeclareInputText{213}{\textquoteright}
968 \DeclareInputMath{214}{\div}
969 \DeclareInputText{215}{\quotedblbase}

970 %

971 \DeclareInputText{216}{\CYRUSHRT}
972 \DeclareInputText{217}{\cyrushrt}
973 ⟨maccyrillic | macukrainian⟩\DeclareInputText{218}{\CYRDZHE}
974 ⟨maccyrillic | macukrainian⟩\DeclareInputText{219}{\cyrdzhe}
975 ⟨pt254⟩\DeclareInputText{218}{\CYRCHVCRS}
976 ⟨pt254⟩\DeclareInputText{219}{\cyrchvcrs}
977 \DeclareInputText{220}{\textnumero}

978 %

979 \DeclareInputText{221}{\CYRYO}

 980 \DeclareInputText{222}{\cyryo}
 981 \DeclareInputText{223}{\cyrya}
 982 \DeclareInputText{224}{\cyra}
 983 \DeclareInputText{225}{\cyrb}
 984 \DeclareInputText{226}{\cyrv}
 985 \DeclareInputText{227}{\cyrg}
 986 \DeclareInputText{228}{\cyrd}
 987 \DeclareInputText{229}{\cyre}
 988 \DeclareInputText{230}{\cyrzh}
 989 \DeclareInputText{231}{\cyrz}
 990 \DeclareInputText{232}{\cyri}
 991 \DeclareInputText{233}{\cyrishrt}
 992 \DeclareInputText{234}{\cyrk}
 993 \DeclareInputText{235}{\cyrl}
 994 \DeclareInputText{236}{\cyrm}
 995 \DeclareInputText{237}{\cyrn}
 996 \DeclareInputText{238}{\cyro}
 997 \DeclareInputText{239}{\cyrp}
 998 \DeclareInputText{240}{\cyrr}
 999 \DeclareInputText{241}{\cyrs}
1000 \DeclareInputText{242}{\cyrt}
1001 \DeclareInputText{243}{\cyru}
1002 \DeclareInputText{244}{\cyrf}
1003 \DeclareInputText{245}{\cyrh}
1004 \DeclareInputText{246}{\cyrc}
1005 \DeclareInputText{247}{\cyrch}
1006 \DeclareInputText{248}{\cyrsh}
1007 \DeclareInputText{249}{\cyrshch}
1008 \DeclareInputText{250}{\cyrhrdsn}
1009 \DeclareInputText{251}{\cyrery}
1010 \DeclareInputText{252}{\cyrsftsn}
1011 \DeclareInputText{253}{\cyrerev}
1012 \DeclareInputText{254}{\cyryu}
1013 ⟨maccyrillic | macukrainian⟩\DeclareInputText{255}{\textcurrency}
1014 ⟨pt254⟩\DeclareInputText{255}{\CYROTLD}

1015 ⟨/maccyrillic | macukrainian | pt254⟩

1.9 Bulgarian MIK (BDS) codepage

It is an MS-DOS codepage used in Bulgaria. This codepage was provided by Georgi Boshnakov and Guentcho Skordev.

1016 ⟨∗MIK⟩

1017 \DeclareInputText{128}{\CYRA}
1018 \DeclareInputText{129}{\CYRB}
1019 \DeclareInputText{130}{\CYRV}
1020 \DeclareInputText{131}{\CYRG}
1021 \DeclareInputText{132}{\CYRD}
1022 \DeclareInputText{133}{\CYRE}
1023 \DeclareInputText{134}{\CYRZH}
1024 \DeclareInputText{135}{\CYRZ}
1025 \DeclareInputText{136}{\CYRI}
1026 \DeclareInputText{137}{\CYRISHRT}
1027 \DeclareInputText{138}{\CYRK}

1028 \DeclareInputText{139}{\CYRL}
1029 \DeclareInputText{140}{\CYRM}
1030 \DeclareInputText{141}{\CYRN}
1031 \DeclareInputText{142}{\CYRO}
1032 \DeclareInputText{143}{\CYRP}
1033 \DeclareInputText{144}{\CYRR}
1034 \DeclareInputText{145}{\CYRS}
1035 \DeclareInputText{146}{\CYRT}
1036 \DeclareInputText{147}{\CYRU}
1037 \DeclareInputText{148}{\CYRF}
1038 \DeclareInputText{149}{\CYRH}
1039 \DeclareInputText{150}{\CYRC}
1040 \DeclareInputText{151}{\CYRCH}
1041 \DeclareInputText{152}{\CYRSH}
1042 \DeclareInputText{153}{\CYRSHCH}
1043 \DeclareInputText{154}{\CYRHRDSN}
1044 \DeclareInputText{155}{\CYRERY}
1045 \DeclareInputText{156}{\CYRSFTSN}
1046 \DeclareInputText{157}{\CYREREV}
1047 \DeclareInputText{158}{\CYRYU}
1048 \DeclareInputText{159}{\CYRYA}
1049 \DeclareInputText{160}{\cyra}
1050 \DeclareInputText{161}{\cyrb}
1051 \DeclareInputText{162}{\cyrv}
1052 \DeclareInputText{163}{\cyrg}
1053 \DeclareInputText{164}{\cyrd}
1054 \DeclareInputText{165}{\cyre}
1055 \DeclareInputText{166}{\cyrzh}
1056 \DeclareInputText{167}{\cyrz}
1057 \DeclareInputText{168}{\cyri}
1058 \DeclareInputText{169}{\cyrishrt}
1059 \DeclareInputText{170}{\cyrk}
1060 \DeclareInputText{171}{\cyrl}
1061 \DeclareInputText{172}{\cyrm}
1062 \DeclareInputText{173}{\cyrn}
1063 \DeclareInputText{174}{\cyro}
1064 \DeclareInputText{175}{\cyrp}
1065 \DeclareInputText{176}{\cyrr}
1066 \DeclareInputText{177}{\cyrs}
1067 \DeclareInputText{178}{\cyrt}
1068 \DeclareInputText{179}{\cyru}
1069 \DeclareInputText{180}{\cyrf}
1070 \DeclareInputText{181}{\cyrh}
1071 \DeclareInputText{182}{\cyrc}
1072 \DeclareInputText{183}{\cyrch}
1073 \DeclareInputText{184}{\cyrsh}
1074 \DeclareInputText{185}{\cyrshch}
1075 \DeclareInputText{186}{\cyrhrdsn}
1076 \DeclareInputText{187}{\cyrery}
1077 \DeclareInputText{188}{\cyrsftsn}
1078 \DeclareInputText{189}{\cyrerev}
1079 \DeclareInputText{190}{\cyryu}
1080 \DeclareInputText{191}{\cyrya}
1081 \DeclareInputText{213}{\textnumero}

1082 \DeclareInputText{214}{\S}
1083 \DeclareInputMath{224}{\alpha}
1084 \DeclareInputMath{225}{\beta}
1085 \DeclareInputMath{226}{\Gamma}
1086 \DeclareInputMath{227}{\pi}
1087 \DeclareInputMath{228}{\Sigma}
1088 \DeclareInputMath{229}{\sigma}
1089 \DeclareInputMath{230}{\mu}
1090 \DeclareInputMath{231}{\tau}
1091 \DeclareInputMath{232}{\Phi}
1092 \DeclareInputMath{233}{\Theta}
1093 \DeclareInputMath{234}{\Omega}
1094 \DeclareInputMath{235}{\delta}
1095 \DeclareInputMath{236}{\infty}
1096 \DeclareInputMath{237}{\emptyset}
1097 \DeclareInputMath{238}{\in}
1098 \DeclareInputMath{239}{\cap}
1099 \DeclareInputMath{240}{\equiv}
1100 \DeclareInputMath{241}{\pm}
1101 \DeclareInputMath{242}{\geq}
1102 \DeclareInputMath{243}{\leq}
1103 \DeclareInputMath{246}{\div}
1104 \DeclareInputMath{247}{\sim}
1105 \DeclareInputText{248}{\textdegree}
1106 \DeclareInputText{249}{\textbullet}
1107 \DeclareInputText{250}{\textperiodcentered}
1108 \DeclareInputMath{251}{\surd}
1109 \DeclareInputMath{252}{\mathnsuperior}
1110 \DeclareInputMath{253}{\mathtwosuperior}
1111 \DeclareInputText{254}{\textblacksquare}
1112 \DeclareInputText{255}{\nobreakspace}

1113 ⟨/MIK⟩

1.10 Mongolian codepages

These codepages were taken from Oliver Corff’s ‘MonTEX’ package (available at CTAN:language/mongolian/montex). Since T2 encodings support the Mongolian Cyrillic script, it is convenient to have support for Mongolian input encodings as well. Pointers to documentation for these codepages are highly appreciated.

1.10.1 CTT Mongolian codepage

1114 ⟨∗CTT⟩

1115 \DeclareInputText{171}{\guillemotleft}
1116 \DeclareInputText{187}{\guillemotright}
1117 \DeclareInputText{192}{\CYRA}
1118 \DeclareInputText{193}{\CYRB}
1119 \DeclareInputText{194}{\CYRV}
1120 \DeclareInputText{195}{\CYRG}
1121 \DeclareInputText{196}{\CYRD}
1122 \DeclareInputText{197}{\CYRE}
1123 \DeclareInputText{168}{\CYRYO}
1124 \DeclareInputText{198}{\CYRZH}
1125 \DeclareInputText{199}{\CYRZ}

1126 \DeclareInputText{200}{\CYRI}
1127 \DeclareInputText{201}{\CYRISHRT}
1128 \DeclareInputText{202}{\CYRK}
1129 \DeclareInputText{203}{\CYRL}
1130 \DeclareInputText{204}{\CYRM}
1131 \DeclareInputText{205}{\CYRN}
1132 \DeclareInputText{206}{\CYRO}
1133 \DeclareInputText{170}{\CYROTLD}
1134 \DeclareInputText{207}{\CYRP}
1135 \DeclareInputText{208}{\CYRR}
1136 \DeclareInputText{209}{\CYRS}
1137 \DeclareInputText{210}{\CYRT}
1138 \DeclareInputText{211}{\CYRU}
1139 \DeclareInputText{175}{\CYRY}
1140 \DeclareInputText{212}{\CYRF}
1141 \DeclareInputText{213}{\CYRH}
1142 \DeclareInputText{214}{\CYRC}
1143 \DeclareInputText{215}{\CYRCH}
1144 \DeclareInputText{216}{\CYRSH}
1145 \DeclareInputText{217}{\CYRSHCH}
1146 \DeclareInputText{218}{\CYRHRDSN}
1147 \DeclareInputText{219}{\CYRERY}
1148 \DeclareInputText{220}{\CYRSFTSN}
1149 \DeclareInputText{221}{\CYREREV}
1150 \DeclareInputText{222}{\CYRYU}
1151 \DeclareInputText{223}{\CYRYA}
1152 \DeclareInputText{224}{\cyra}
1153 \DeclareInputText{225}{\cyrb}
1154 \DeclareInputText{226}{\cyrv}
1155 \DeclareInputText{227}{\cyrg}
1156 \DeclareInputText{228}{\cyrd}
1157 \DeclareInputText{229}{\cyre}
1158 \DeclareInputText{184}{\cyryo}
1159 \DeclareInputText{230}{\cyrzh}
1160 \DeclareInputText{231}{\cyrz}
1161 \DeclareInputText{232}{\cyri}
1162 \DeclareInputText{233}{\cyrishrt}
1163 \DeclareInputText{234}{\cyrk}
1164 \DeclareInputText{235}{\cyrl}
1165 \DeclareInputText{236}{\cyrm}
1166 \DeclareInputText{237}{\cyrn}
1167 \DeclareInputText{238}{\cyro}
1168 \DeclareInputText{186}{\cyrotld}
1169 \DeclareInputText{239}{\cyrp}
1170 \DeclareInputText{240}{\cyrr}
1171 \DeclareInputText{241}{\cyrs}
1172 \DeclareInputText{242}{\cyrt}
1173 \DeclareInputText{243}{\cyru}
1174 \DeclareInputText{191}{\cyry}
1175 \DeclareInputText{244}{\cyrf}
1176 \DeclareInputText{245}{\cyrh}
1177 \DeclareInputText{246}{\cyrc}
1178 \DeclareInputText{247}{\cyrch}
1179 \DeclareInputText{248}{\cyrsh}

1180 \DeclareInputText{249}{\cyrshch}
1181 \DeclareInputText{250}{\cyrhrdsn}
1182 \DeclareInputText{251}{\cyrery}
1183 \DeclareInputText{252}{\cyrsftsn}
1184 \DeclareInputText{253}{\cyrerev}
1185 \DeclareInputText{254}{\cyryu}
1186 \DeclareInputText{255}{\cyrya}

1187 ⟨/CTT⟩

1.10.2 DBK Mongolian codepage

1188 ⟨∗DBK⟩

1189 \DeclareInputText{128}{\CYRA}
1190 \DeclareInputText{129}{\CYRB}
1191 \DeclareInputText{130}{\CYRV}
1192 \DeclareInputText{131}{\CYRG}
1193 \DeclareInputText{132}{\CYRD}
1194 \DeclareInputText{133}{\CYRE}
1195 \DeclareInputText{134}{\CYRYO}
1196 \DeclareInputText{135}{\CYRZH}
1197 \DeclareInputText{136}{\CYRZ}
1198 \DeclareInputText{137}{\CYRI}
1199 \DeclareInputText{139}{\CYRISHRT}
1200 \DeclareInputText{140}{\CYRK}
1201 \DeclareInputText{142}{\CYRL}
1202 \DeclareInputText{143}{\CYRM}
1203 \DeclareInputText{144}{\CYRN}
1204 \DeclareInputText{145}{\CYRO}
1205 \DeclareInputText{146}{\CYROTLD}
1206 \DeclareInputText{147}{\CYRP}
1207 \DeclareInputText{148}{\CYRR}
1208 \DeclareInputText{149}{\CYRS}
1209 \DeclareInputText{150}{\CYRT}
1210 \DeclareInputText{151}{\CYRU}
1211 \DeclareInputText{152}{\CYRY}
1212 \DeclareInputText{153}{\CYRF}
1213 \DeclareInputText{154}{\CYRH}
1214 \DeclareInputText{155}{\CYRC}
1215 \DeclareInputText{156}{\CYRCH}
1216 \DeclareInputText{157}{\CYRSH}
1217 \DeclareInputText{158}{\CYRSHCH}
1218 \DeclareInputText{159}{\CYRHRDSN}
1219 \DeclareInputText{160}{\CYRERY}
1220 \DeclareInputText{161}{\CYRSFTSN}
1221 \DeclareInputText{162}{\CYREREV}
1222 \DeclareInputText{163}{\CYRYU}
1223 \DeclareInputText{164}{\CYRYA}
1224 \DeclareInputText{165}{\cyra}
1225 \DeclareInputText{166}{\cyrb}
1226 \DeclareInputText{167}{\cyrv}
1227 \DeclareInputText{168}{\cyrg}
1228 \DeclareInputText{169}{\cyrd}
1229 \DeclareInputText{170}{\cyre}
1230 \DeclareInputText{171}{\cyryo}
1231 \DeclareInputText{172}{\cyrzh}

1232 \DeclareInputText{173}{\cyrz}
1233 \DeclareInputText{174}{\cyri}
1234 \DeclareInputText{175}{\cyrishrt}
1235 \DeclareInputText{225}{\cyrk}
1236 \DeclareInputText{226}{\cyrl}
1237 \DeclareInputText{227}{\cyrm}
1238 \DeclareInputText{228}{\cyrn}
1239 \DeclareInputText{229}{\cyro}
1240 \DeclareInputText{230}{\cyrotld}
1241 \DeclareInputText{231}{\cyrp}
1242 \DeclareInputText{232}{\cyrr}
1243 \DeclareInputText{233}{\cyrs}
1244 \DeclareInputText{234}{\cyrt}
1245 \DeclareInputText{235}{\cyru}
1246 \DeclareInputText{236}{\cyry}
1247 \DeclareInputText{237}{\cyrf}
1248 \DeclareInputText{238}{\cyrh}
1249 \DeclareInputText{239}{\cyrc}
1250 \DeclareInputText{241}{\cyrch}
1251 \DeclareInputText{242}{\cyrsh}
1252 \DeclareInputText{243}{\cyrshch}
1253 \DeclareInputText{244}{\cyrhrdsn}
1254 \DeclareInputText{245}{\cyrery}
1255 \DeclareInputText{246}{\cyrsftsn}
1256 \DeclareInputText{247}{\cyrerev}
1257 \DeclareInputText{248}{\cyryu}
1258 \DeclareInputText{249}{\cyrya}

1259 ⟨/DBK⟩

1.10.3 MNK Mongolian codepage

1260 ⟨∗MNK⟩

1261 \DeclareInputText{128}{\CYRA}
1262 \DeclareInputText{129}{\CYRB}
1263 \DeclareInputText{130}{\CYRV}
1264 \DeclareInputText{131}{\CYRG}
1265 \DeclareInputText{132}{\CYRD}
1266 \DeclareInputText{133}{\CYRE}
1267 \DeclareInputText{134}{\CYRYO}
1268 \DeclareInputText{135}{\CYRZH}
1269 \DeclareInputText{136}{\CYRZ}
1270 \DeclareInputText{137}{\CYRI}
1271 \DeclareInputText{138}{\CYRISHRT}
1272 \DeclareInputText{139}{\CYRK}
1273 \DeclareInputText{140}{\CYRL}
1274 \DeclareInputText{141}{\CYRM}
1275 \DeclareInputText{142}{\CYRN}
1276 \DeclareInputText{143}{\CYRO}
1277 \DeclareInputText{144}{\CYROTLD}
1278 \DeclareInputText{145}{\CYRP}
1279 \DeclareInputText{146}{\CYRR}
1280 \DeclareInputText{147}{\CYRS}
1281 \DeclareInputText{148}{\CYRT}
1282 \DeclareInputText{149}{\CYRU}
1283 \DeclareInputText{150}{\CYRY}

1284 \DeclareInputText{151}{\CYRF}
1285 \DeclareInputText{152}{\CYRH}
1286 \DeclareInputText{153}{\CYRC}
1287 \DeclareInputText{154}{\CYRCH}
1288 \DeclareInputText{155}{\CYRSH}
1289 \DeclareInputText{156}{\CYRSHCH}
1290 \DeclareInputText{157}{\CYRHRDSN}
1291 \DeclareInputText{158}{\CYRERY}
1292 \DeclareInputText{159}{\CYRSFTSN}
1293 \DeclareInputText{160}{\CYREREV}
1294 \DeclareInputText{161}{\CYRYU}
1295 \DeclareInputText{162}{\CYRYA}
1296 \DeclareInputText{163}{\cyra}
1297 \DeclareInputText{164}{\cyrb}
1298 \DeclareInputText{165}{\cyrv}
1299 \DeclareInputText{166}{\cyrg}
1300 \DeclareInputText{167}{\cyrd}
1301 \DeclareInputText{168}{\cyre}
1302 \DeclareInputText{169}{\cyryo}
1303 \DeclareInputText{170}{\cyrzh}
1304 \DeclareInputText{173}{\cyrz}
1305 \DeclareInputText{224}{\cyri}
1306 \DeclareInputText{225}{\cyrishrt}
1307 \DeclareInputText{226}{\cyrk}
1308 \DeclareInputText{227}{\cyrl}
1309 \DeclareInputText{228}{\cyrm}
1310 \DeclareInputText{229}{\cyrn}
1311 \DeclareInputText{230}{\cyro}
1312 \DeclareInputText{231}{\cyrotld}
1313 \DeclareInputText{232}{\cyrp}
1314 \DeclareInputText{233}{\cyrr}
1315 \DeclareInputText{234}{\cyrs}
1316 \DeclareInputText{235}{\cyrt}
1317 \DeclareInputText{236}{\cyru}
1318 \DeclareInputText{237}{\cyry}
1319 \DeclareInputText{238}{\cyrf}
1320 \DeclareInputText{239}{\cyrh}
1321 \DeclareInputText{240}{\cyrc}
1322 \DeclareInputText{241}{\cyrch}
1323 \DeclareInputText{242}{\cyrsh}
1324 \DeclareInputText{243}{\cyrshch}
1325 \DeclareInputText{244}{\cyrhrdsn}
1326 \DeclareInputText{245}{\cyrery}
1327 \DeclareInputText{248}{\cyrsftsn}
1328 \DeclareInputText{252}{\cyrerev}
1329 \DeclareInputText{253}{\cyryu}
1330 \DeclareInputText{254}{\cyrya}

1331 ⟨/MNK⟩

1.10.4 MOS Mongolian codepage

1332 ⟨∗MOS⟩

1333 \DeclareInputText{128}{\CYRA}
1334 \DeclareInputText{129}{\CYRB}
1335 \DeclareInputText{130}{\CYRV}

1336 \DeclareInputText{131}{\CYRG}
1337 \DeclareInputText{132}{\CYRD}
1338 \DeclareInputText{133}{\CYRE}
1339 \DeclareInputText{160}{\CYRYO}
1340 \DeclareInputText{134}{\CYRZH}
1341 \DeclareInputText{135}{\CYRZ}
1342 \DeclareInputText{136}{\CYRI}
1343 \DeclareInputText{137}{\CYRISHRT}
1344 \DeclareInputText{138}{\CYRK}
1345 \DeclareInputText{139}{\CYRL}
1346 \DeclareInputText{140}{\CYRM}
1347 \DeclareInputText{141}{\CYRN}
1348 \DeclareInputText{142}{\CYRO}
1349 \DeclareInputText{153}{\CYROTLD}
1350 \DeclareInputText{143}{\CYRP}
1351 \DeclareInputText{144}{\CYRR}
1352 \DeclareInputText{145}{\CYRS}
1353 \DeclareInputText{146}{\CYRT}
1354 \DeclareInputText{147}{\CYRU}
1355 \DeclareInputText{154}{\CYRY}
1356 \DeclareInputText{148}{\CYRF}
1357 \DeclareInputText{149}{\CYRH}
1358 \DeclareInputText{150}{\CYRC}
1359 \DeclareInputText{151}{\CYRCH}
1360 \DeclareInputText{152}{\CYRSH}
1361 \DeclareInputText{164}{\CYRSHCH}
1362 \DeclareInputText{162}{\CYRHRDSN}
1363 \DeclareInputText{155}{\CYRERY}
1364 \DeclareInputText{156}{\CYRSFTSN}
1365 \DeclareInputText{157}{\CYREREV}
1366 \DeclareInputText{158}{\CYRYU}
1367 \DeclareInputText{159}{\CYRYA}
1368 \DeclareInputText{224}{\cyra}
1369 \DeclareInputText{225}{\cyrb}
1370 \DeclareInputText{226}{\cyrv}
1371 \DeclareInputText{227}{\cyrg}
1372 \DeclareInputText{228}{\cyrd}
1373 \DeclareInputText{229}{\cyre}
1374 \DeclareInputText{161}{\cyryo}
1375 \DeclareInputText{230}{\cyrzh}
1376 \DeclareInputText{231}{\cyrz}
1377 \DeclareInputText{232}{\cyri}
1378 \DeclareInputText{233}{\cyrishrt}
1379 \DeclareInputText{234}{\cyrk}
1380 \DeclareInputText{235}{\cyrl}
1381 \DeclareInputText{236}{\cyrm}
1382 \DeclareInputText{237}{\cyrn}
1383 \DeclareInputText{238}{\cyro}
1384 \DeclareInputText{249}{\cyrotld}
1385 \DeclareInputText{239}{\cyrp}
1386 \DeclareInputText{240}{\cyrr}
1387 \DeclareInputText{241}{\cyrs}
1388 \DeclareInputText{242}{\cyrt}
1389 \DeclareInputText{243}{\cyru}

1390 \DeclareInputText{250}{\cyry}
1391 \DeclareInputText{244}{\cyrf}
1392 \DeclareInputText{245}{\cyrh}
1393 \DeclareInputText{246}{\cyrc}
1394 \DeclareInputText{247}{\cyrch}
1395 \DeclareInputText{248}{\cyrsh}
1396 \DeclareInputText{165}{\cyrshch}
1397 \DeclareInputText{163}{\cyrhrdsn}
1398 \DeclareInputText{251}{\cyrery}
1399 \DeclareInputText{252}{\cyrsftsn}
1400 \DeclareInputText{253}{\cyrerev}
1401 \DeclareInputText{254}{\cyryu}
1402 \DeclareInputText{168}{\cyrya}

1403 ⟨/MOS⟩

1.10.5 NCC Mongolian codepage

1404 ⟨∗NCC⟩

1405 \DeclareInputText{128}{\CYRA}
1406 \DeclareInputText{129}{\CYRB}
1407 \DeclareInputText{130}{\CYRV}
1408 \DeclareInputText{131}{\CYRG}
1409 \DeclareInputText{132}{\CYRD}
1410 \DeclareInputText{133}{\CYRE}
1411 \DeclareInputText{134}{\CYRYO}
1412 \DeclareInputText{135}{\CYRZH}
1413 \DeclareInputText{136}{\CYRZ}
1414 \DeclareInputText{137}{\CYRI}
1415 \DeclareInputText{139}{\CYRISHRT}
1416 \DeclareInputText{140}{\CYRK}
1417 \DeclareInputText{142}{\CYRL}
1418 \DeclareInputText{143}{\CYRM}
1419 \DeclareInputText{144}{\CYRN}
1420 \DeclareInputText{145}{\CYRO}
1421 \DeclareInputText{146}{\CYROTLD}
1422 \DeclareInputText{147}{\CYRP}
1423 \DeclareInputText{148}{\CYRR}
1424 \DeclareInputText{149}{\CYRS}
1425 \DeclareInputText{150}{\CYRT}
1426 \DeclareInputText{151}{\CYRU}
1427 \DeclareInputText{152}{\CYRY}
1428 \DeclareInputText{153}{\CYRF}
1429 \DeclareInputText{154}{\CYRH}
1430 \DeclareInputText{155}{\CYRC}
1431 \DeclareInputText{156}{\CYRCH}
1432 \DeclareInputText{157}{\CYRSH}
1433 \DeclareInputText{158}{\CYRSHCH}
1434 \DeclareInputText{159}{\CYRHRDSN}
1435 \DeclareInputText{160}{\CYRERY}
1436 \DeclareInputText{161}{\CYRSFTSN}
1437 \DeclareInputText{162}{\CYREREV}
1438 \DeclareInputText{163}{\CYRYU}
1439 \DeclareInputText{164}{\CYRYA}
1440 \DeclareInputText{165}{\cyra}
1441 \DeclareInputText{166}{\cyrb}

1442 \DeclareInputText{167}{\cyrv}
1443 \DeclareInputText{168}{\cyrg}
1444 \DeclareInputText{169}{\cyrd}
1445 \DeclareInputText{170}{\cyre}
1446 \DeclareInputText{171}{\cyryo}
1447 \DeclareInputText{172}{\cyrzh}
1448 \DeclareInputText{173}{\cyrz}
1449 \DeclareInputText{225}{\cyri}
1450 \DeclareInputText{226}{\cyrishrt}
1451 \DeclareInputText{227}{\cyrk}
1452 \DeclareInputText{228}{\cyrl}
1453 \DeclareInputText{229}{\cyrm}
1454 \DeclareInputText{230}{\cyrn}
1455 \DeclareInputText{231}{\cyro}
1456 \DeclareInputText{232}{\cyrotld}
1457 \DeclareInputText{233}{\cyrp}
1458 \DeclareInputText{234}{\cyrr}
1459 \DeclareInputText{235}{\cyrs}
1460 \DeclareInputText{236}{\cyrt}
1461 \DeclareInputText{237}{\cyru}
1462 \DeclareInputText{238}{\cyry}
1463 \DeclareInputText{239}{\cyrf}
1464 \DeclareInputText{240}{\cyrh}
1465 \DeclareInputText{241}{\cyrc}
1466 \DeclareInputText{242}{\cyrch}
1467 \DeclareInputText{243}{\cyrsh}
1468 \DeclareInputText{244}{\cyrshch}
1469 \DeclareInputText{245}{\cyrhrdsn}
1470 \DeclareInputText{246}{\cyrery}
1471 \DeclareInputText{247}{\cyrsftsn}
1472 \DeclareInputText{248}{\cyrerev}
1473 \DeclareInputText{249}{\cyryu}
1474 \DeclareInputText{251}{\cyrya}

1475 ⟨/NCC⟩

1.10.6 MLS Mongolian codepage

1476 ⟨∗MLS⟩

1477 \DeclareInputText{128}{\CYRB}
1478 \DeclareInputText{129}{\cyry}
1479 \DeclareInputText{130}{\CYRD}
1480 \DeclareInputText{131}{\CYRYO}
1481 \DeclareInputText{132}{\cyrerev}
1482 \DeclareInputText{133}{\CYRZH}
1483 \DeclareInputText{134}{\CYRZ}
1484 \DeclareInputText{135}{\CYRI}
1485 \DeclareInputText{136}{\CYRISHRT}
1486 \DeclareInputText{137}{\cyryo}
1487 \DeclareInputText{138}{\CYRL}
1488 \DeclareInputText{139}{\cyrishrt}
1489 \DeclareInputText{140}{\CYROTLD}
1490 \DeclareInputText{141}{\CYRP}
1491 \DeclareInputText{142}{\CYREREV}
1492 \DeclareInputText{143}{\CYRU}
1493 \DeclareInputText{144}{\CYRF}

1494 \DeclareInputText{145}{\CYRC}
1495 \DeclareInputText{146}{\CYRCH}
1496 \DeclareInputText{147}{\CYRSH}
1497 \DeclareInputText{148}{\cyrotld}
1498 \DeclareInputText{149}{\CYRSHCH}
1499 \DeclareInputText{150}{\CYRHRDSN}
1500 \DeclareInputText{151}{\CYRERY}
1501 \DeclareInputText{152}{\CYRSFTSN}
1502 \DeclareInputText{153}{\CYROTLD}
1503 \DeclareInputText{154}{\CYRY}
1504 \DeclareInputText{155}{\CYREREV}
1505 \DeclareInputText{156}{\CYRYU}
1506 \DeclareInputText{157}{\CYRYA}
1507 \DeclareInputText{158}{\cyrb}
1508 \DeclareInputText{159}{\cyrv}
1509 \DeclareInputText{160}{\cyrg}
1510 \DeclareInputText{161}{\cyrd}
1511 \DeclareInputText{162}{\cyrzh}
1512 \DeclareInputText{163}{\cyrz}
1513 \DeclareInputText{164}{\cyri}
1514 \DeclareInputText{165}{\cyrishrt}
1515 \DeclareInputText{166}{\cyrk}
1516 \DeclareInputText{167}{\cyrl}
1517 \DeclareInputText{168}{\cyrm}
1518 \DeclareInputText{169}{\cyrn}
1519 \DeclareInputText{170}{\cyrotld}
1520 \DeclareInputText{171}{\cyrp}
1521 \DeclareInputText{172}{\cyrt}
1522 \DeclareInputText{173}{\cyry}
1523 \DeclareInputText{174}{\guillemotleft}
1524 \DeclareInputText{175}{\guillemotright}
1525 \DeclareInputText{176}{\cyrf}
1526 \DeclareInputText{177}{\cyrc}
1527 \DeclareInputText{178}{\cyrch}
1528 \DeclareInputText{180}{\cyrsh}
1529 \DeclareInputText{181}{\cyrshch}
1530 \DeclareInputText{182}{\cyrhrdsn}
1531 \DeclareInputText{183}{\cyrery}
1532 \DeclareInputText{184}{\cyrsftsn}
1533 \DeclareInputText{189}{\cyrerev}
1534 \DeclareInputText{190}{\cyryu}
1535 \DeclareInputText{193}{\cyrya}
1536 \DeclareInputText{226}{\CYRG}
1537 \DeclareInputText{225}{\ss}
1538 \DeclareInputText{231}{\ensuremath{\gamma}}
1539 \DeclareInputText{255}{\nobreakspace}

Bicig Letters. These are traditional (non-Cyrillic) Mongolian letters, which are not supported by Cyrillic T2 encodings. To use these letters you should install the LMS font encoding definition file and Mongolian fonts contained in the MonTEX package. These letters coexist with Cyrillic in one input encoding.

1540 \DeclareInputText{194}{\titem}
1541 \DeclareInputText{195}{\shud}
1542 \DeclareInputText{197}{\secondaryshud}

1543 \DeclareInputText{198}{\shilbe}
1544 \DeclareInputText{199}{\gedes}
1545 \DeclareInputText{207}{\secondarygedes}
1546 \DeclareInputText{208}{\cegteishud}
1547 \DeclareInputText{209}{\lewer}
1548 \DeclareInputText{210}{\suuliinlewer}
1549 \DeclareInputText{211}{\tertiarylewer}
1550 \DeclareInputText{212}{\mewer}
1551 \DeclareInputText{213}{\suuliinmewer}
1552 \DeclareInputText{214}{\xewteeqix}
1553 \DeclareInputText{215}{\dawxarcegtxewteeqix}
1554 \DeclareInputText{216}{\halfnum}
1555 \DeclareInputText{219}{\num}
1556 \DeclareInputText{220}{\halfnumtgedes}
1557 \DeclareInputText{221}{\numtaigedes}
1558 \DeclareInputText{222}{\buruuxarsangedes}
1559 \DeclareInputText{223}{\gedesteishilbe}
1560 \DeclareInputText{224}{\erweeljinshilbe}
1561 \DeclareInputText{227}{\secerweeljin}
1562 \DeclareInputText{228}{\bosooshilbe}
1563 \DeclareInputText{229}{\etgershilbe}
1564 \DeclareInputText{230}{\zawj}
1565 \DeclareInputText{232}{\suuliinzawj}
1566 \DeclareInputText{233}{\dawxarcegtzawj}
1567 \DeclareInputText{234}{\sereeewer}
1568 \DeclareInputText{235}{\matgarshilbe}
1569 \DeclareInputText{236}{\bituushilbe}
1570 \DeclareInputText{237}{\secondaryqagt}
1571 \DeclareInputText{238}{\qagt}
1572 \DeclareInputText{239}{\secnumtdelbenqix}
1573 \DeclareInputText{240}{\numtdelbenqix}
1574 \DeclareInputText{241}{\secsertenqixtnum}
1575 \DeclareInputText{242}{\sertenqixtnum}
1576 \DeclareInputText{243}{\zadgaizardigt}
1577 \DeclareInputText{244}{\bituuzardigt}
1578 \DeclareInputText{245}{\malgaitaititem}
1579 \DeclareInputText{246}{\suul}
1580 \DeclareInputText{247}{\orxic}
1581 \DeclareInputText{248}{\biodoisuul}
1582 \DeclareInputText{249}{\bagodoisuul}
1583 \DeclareInputText{250}{\nceg}
1584 \DeclareInputText{251}{\gceg}
1585 \DeclareInputText{252}{\ceg}
1586 \DeclareInputText{253}{\dorwoljin}

1587 ⟨/MLS⟩

Finally, we reset the category code of the at sign at the end of all .def files.

1588 \makeatother